OpenStudy (anonymous):

If f(x) = 3x – 2 and g(x) = 6 – 4x, find f(x) + g(x). A. 7 + 4x B. 4 – x C. –x – 4 D. –4 – 7x what is there answer please?

OpenStudy (anonymous):

If f(x) = 4x – 5 and g(x) = 3 – 2x, find f(x) + g(x). A. 2x – 2 B. 6x – 2 C. 2x – 8 D. 6x – 8

OpenStudy (paki):

simple... add both the functions... and combine the like terms... and solve then...

OpenStudy (anonymous):

If f(x) = 4x – 7 and g(x) = 2x + 4, evaluate f(x) + g(x) for x = –3. A. –52 B. –21 C. –22 D. –15